OpenStudy (yada):

y=10000*0.8^x if y is dollas and x is year. please help

OpenStudy (anonymous):

\[y = 10000 * .8^x\]

OpenStudy (yada):

yes. I could not solve that question

OpenStudy (anonymous):

what are you trying to solve?

OpenStudy (yada):

i could fine in percent/year

OpenStudy (yada):

if y is money and x is year so hur big is cost in percent per year

OpenStudy (anonymous):

@across

OpenStudy (across):

I am so confused. What are you trying to do? Solve for \(x\)?

OpenStudy (yada):

yes

OpenStudy (across):

Do you know how to work with logarithms?

OpenStudy (across):

\[ y=10000\cdot0.8^x\\ \log_{0.8}y=\log_{0.8}10000+\log_{0.8}0.8^x\\ \log_{0.8}y=\log_{0.8}10000+x\\ x=\log_{0.8}y-\log_{0.8}10000\\ x=\log_{0.8}\frac{y}{10000} \]

OpenStudy (yada):

i do not know that. thats why i could not do that question also its not in the book :((

OpenStudy (yada):

how to get the percent/year because i could not know the result of y